# Large Language Models and Games: A Survey and Roadmap

### v2

@@ 3.5 Analyst
So far, language models (both large and small) have been used for clustering player behaviors. Player2vec [66] trained the Longformer transformer architecture [67], with up to 121 million parameters, on a corpus of game events stored in JSON format which captured player interactions with a casual mobile game. Using dimensionality reduction methods on the latent vectors of the last layer of a pre-trained transformer, the authors discovered eight clusters with player traits that could be useful for market research, such as “lean-in casual economy aware”. In such work, the strength of LLMs is their ability to process structured or unstructured data without the need for extensive data pre-processing. However, as with previous experiments in bottom-up clustering of players [68, 69], the resulting clusters need to be interpreted by expert game analysts (or the game’s designers) to derive some meaningful player types. In this form, current research points to LLMs acting as analyst assistants rather than independent analysts; we revisit future directions in this vein in Section 4.
Representing game logs as text interpreted by an LLM need not be the end-goal for analysis. The LLM representations of game logs can also be used to find common patterns that can be used for other purposes such as gameplay (or gameplay footage) similarity. Indicatively, Trivedi et al. [70] used LLMs on recorded gameplay action logs to establish action similarity. This similarity was in turn applied to train video encodings on gameplay footage of these recorded actions, in order to align the visual latent vectors to the action logs. While the work focused on the computer vision task of representation learning for gameplay footage, better (learned) representations of gameplay pixels would be more generalizable across games in downstream tasks including action recognition or affect modeling; we expand on these downstream tasks in Section 4.
@@ 3.6 Game Master
A Game Master (GM) in tabletop role-playing games (TTRPGs) is the person who creates the plot of a game, its characters, and narrative. GMs wear many hats during the course of the game session [71]; they prepare and adapt stories before sessions, guide gameplay during, and follow up with players afterward [72]. Digital games have mostly pre-scripted stories or level progressions, and their players have a restricted range of affordances, compared to TTRPG players whose actions are only limited by their imagination. Similarly, the story told around the table can take any direction. Since human GMs mostly communicate about the gameworld, story, game state and action resolutions via natural language (although props such as maps, miniatures, hand-outs are also common), the potential of LLMs as a GM is often mentioned both in research circles and TTRPG discussion boards. LLMs as GMs also open the potential for solo play, while a TTRPG requires at least one player and a human GM.
@@
One of the first notable text adventures managed by a fine-tuned version of GPT-2 is AI Dungeon [73]. It is an online [^4] https://play.aidungeon.com/ interactive chat-based storytelling application where the player takes actions through semantic input alone. The LLM continues the story based on the player’s input, in the fashion of a human GM. The game has evolved since its creation to make use of more recent LLM models, which the player can choose from before starting a play session. Different game world settings are also offered, and players are also able to share the stories they create. Similar games have emerged online ever since [^5] https://koboldai.net, https://www.hiddendoor.co, and a freely available code repository, Kobold AI Client [^6] https://github.com/KoboldAI/KoboldAI-Client, allows a local or remote installation of a client for such LLM-run games. Some of these games also use Stable Diffusion text-to-image models [74] to generate visuals accompanying different parts of the narrative. More recent work [75] investigates how different characterizations of the GM and their way of presenting events to the players impacts the overall experience in TTRPGs. This study presented an online interface for a custom game based on a Dungeons & Dragons (TSR, 1974) (D&D) fantasy setting. LLMs were ascribed different GM roles through a number of prompts, and human players participated in 45-minute games (split into 3 equal sub-sessions), with an overall positive response to the automated GM.
@@
In lieu of replacing a human GM, LLMs have also been employed as GM assistants. CALYPSO [76] is a set of tools running on a Discord server which the GM can query either to generate random encounters, brainstorm ideas, or alternatively chat with a fictional character in a D&D TTRPG setting. CALYPSO highlights that hallucinations of GPT-3 can have both positive effects when it generates plausible details not included in descriptions published in the original game manual (e.g. the shapes of creatures’ eyes) and also negative effects when the created details are outright incorrect (e.g. describing the wings of a canonically wingless creature). In addition, the model’s preconditioning to avoid racial bias was found to occasionally prevent it from generating racial details of fantasy creatures in the game. Other work used smaller GPT models to improvise in-game conversations [77] by monitoring and transcribing verbal exchanges between the GM and the players, and attempting to generate appropriate responses. This example was integrated into “Shoelace” [78], which is itself a GM assisting tool helping with content lookup by creating a node-based plan of the game narrative and encounters. The versatility of LLMs given their ability to rapidly process text input paves the way for their integration into the multitude of existing tools and aids for human GMs.
@@ 3.7 Game Mechanic
Another strength of LLMs is language synthesis. Huang at al. [82] used GPT-3 to generate new words from the combination of two user-selected words. These are used to progress in a text-based game scenario, where the player wins by unlocking goal words. Similarly, language synthesis is leveraged by Infinite Craft [^8] https://neal.fun/infinite-craft/, an “alchemy” game, in which the player combines elements to produce new ones (see Fig. 3). In Infinite Craft, the player begins with a set of core elements (water, fire, wind and earth). But while the former have a set of interactions defined manually by the designer, Infinite Craft prompts LLaMA 2 [18] to imagine the product of the combination of these elements [83]. Judging from gameplay, it appears that for each distinct combination, LLaMA is prompted to produce the result only once, with the product stored in a database for future reference. Thus seemingly anything in the language model’s vocabulary might “emerge” from the combination of these elements, including all 50 states [^9] https://x.com/FeralFlex/status/1758332430136615298, “Dream” [^10] https://x.com/slutzsmp/status/1760394123243135169, and the fictional “Super Stonedosaurus Tacosaurus Rex” [^11] https://x.com/pvtspicy/status/1759316982984237139. On occasion, the model can choose to return one of the combined elements, or refuse to combine (e.g. very lengthy or complex) elements.
@@ 3.8 Automated Designer
Any PCG method that is trained on available content corpora fits under the Procedural Content Generation via Machine Learning (PCGML) paradigm [21]. Strictly speaking the original PCGML framework of 2018 did not consider LLMs; instead it relied on machine learning methods such as autoencoders and LSTMs. However, important challenges of PCGML remain when considering LLMs for PCG: notably, the reliance on high-quality, machine-readable datasets from human-authored levels. While some datasets exist for arcade game levels [84], for most games the content remains both unavailable and protected by intellectual property (IP) laws. We revisit this issue in Section 6.
Prior work in PCG has demonstrated that tile-based game levels can be reliably generated with sequence-based prediction models (e.g. LSTMs) from a modest set of examples, by treating such levels as linear sequences of tile types in raster order [85, 86].
More recently, this approach has been extended to modern LLMs pre-trained on natural language (instead of sequence models trained from the ground up). Todd et al. [87] fine-tuned a GPT-2 model on a large dataset of Sokoban (Thinking Rabbit, 1982) levels and, at test time, samples from the model to produce novel puzzles (see Figure 4). Interestingly, their results indicate that while the GPT-2 model struggles when the fine-tuning dataset is restricted in size, GPT-3 (and, presumably, larger models released since then) are better able to accommodate limited training sets.
@@
Using a similar approach, MarioGPT, trains a GPT-2 model on a relatively small dataset of Super Mario Bros (Nintendo, 1985) levels [89]. MarioGPT overcomes the issue of data sparsity by using the initial dataset as the starting point for an evolutionary algorithm. Existing levels are selected and then sections of the level are mutated by sampling from the GPT model and then correcting the border between the re-generated section and the rest of the level with a similarly-trained BERT (i.e. bi-directional) model [2]. This approach produces a large and diverse set of playable levels, despite starting from less than 20 levels.
Both aforementioned GPT-based level generation approaches also show the promise of incorporating natural language instructions to produce conditional level generators, either by prefixing game levels in the training dataset with desired level characteristics [87] or by embedding user instructions and allowing the model to attend to the embedding during generation [89]. A recent example is Cardistry [90], which used GPT-3.5 to transform a short personal narrative into a set of playing cards. Along with text information, the LLM creates prompts for DALL-E [91], completing the generation pipeline with playing card art.
User requests however are typically fuzzy and can be misinterpreted by the LLM. Hu et al. [54] first “refine” user requests for the generation of Minecraft structures using an LLM, adding knowledge specific to the domain such as block palette and building dimensions. Then, a second LLM is employed to generate the description of the structure that can be interpreted and placed in the game.
Figure 5: A screenshot of a CrawLLM game instance, generated for the theme of Ancient Egypt [92]. Image used with permission.
@@ 3.9 Design Assistant
An AI for design assistance can provide several benefits to the creative process. Depending on the type of tool, type of AI, and type of creative process, the AI can minimize development time and cost, reduce human effort, support collaboration among members of a design , or elicit a user’s creativity [94]. So far, in games, most of the AI-powered design assistant tools focus on autocompleting a human’s in-progress design [95] or providing many possible suggestions for the designer to consider [96, 97, 98, 99, 100, 101]. Ideally, we would want an LLM that can act as a human colleague that we can bounce off ideas to and collaborate with. Such LLMs are still beyond the current state of the art [102]. Existing tools implement co-creating [103] LLMs at different levels of control, which we can explore under the existing Co-Creative Framework for Interaction Design [104]. Focusing on the interaction over the artifact, the LLM can be of conceptual assistance, providing high-level guidance which is not game-ready. This would require that the designer adapts and curates the AI output in a way that fits their own vision and the constraints of the game. This implies that the LLM contributes very little in the actual artifact generation, merely providing new suggestions. By allowing the LLM to also refine or transform the artifact, we see a procedural assistance by the LLM. Interacting with the user, the LLM can produce increasingly more final versions of the intended artifact. The LLM is expected to understand the context of the game for which the content is intended, in order to provide meaningful assistance. However, the LLM does not need to produce a final, playable artifact but could instead simply provide the next creative step for discussion with the designer [105]. Moreover, the designer is ultimately responsible for curating and adapting the generated content, as well as deciding when the co-creative process is completed [106, 107, 108]. Finally, if the LLM is allowed to directly create and alter the artifact based on user requirements, we say it provides production assistance. This is the closest level to PCG (see Section 3.8), but is different in that the designer remains in control and can refine their specification or reject a created artifact (versus an autonomous generator which directly sends content to the player). As expected, however, the AI operates in a much more constrained space in this scenario as it must account for all other game mechanics (the design of which are presumed finalized) and designer goals which are somehow encoded or presumed via learned designer models [109].
@@
One can argue that existing interfaces with LLMs, FMs, and Large Multimodal Models (LMMs) act as design assistants. The designer provides their specifications and receives one (in LLMs) or multiple (in AI image generators) suggestions that they can further refine. Many creatives report using such interfaces for brainstorming and concept development [110], including game developers [111]. However, the applicability of LLMs as design assistants is somewhat limited, reverting only to conceptual assistance. Similarly, their potential for refining an existing idea (i.e. offering procedural assistance) is underexplored, as we discuss in Section 4.
@@
Conceptual assistance is thus the easiest for LLMs, and is the first case explored in games. Charity et al. [112] envision design assistance as a tool that combines the game description provided by the user with existing knowledge of similar games to suggest possible game features back to the user. The features suggested are fairly generic, few-word guidelines (e.g.: “learn new combat”) which would need extensive design effort and creativity to transform into an implementable and coherent game design. When asking an LLM for specific game features to implement in a digital game, however, players would find them less compelling than human-designed ones [113]. The suggestions, however, were still useful for game designers, as they provided a different perspective that could kickstart their creation process. A thorough analysis of strengths and weaknesses of LLMs in this role can be drawn form Project AVA [114], a non-commercial digital game developed at Keywords Studios with the assistance of LLMs and FMs for multiple aspects of the typical game development pipeline. LLMs are shown to help greatly in giving inspiration to the designers (albeit not being creative themselves), provide simple starting code for the game logic, and even provide base assistance to developers with surfacing errors in the code during development. LLMs however often fall short on anything more involved or requiring further domain knowledge, such as requesting code for specific game logic or understanding functionality in UI elements. Similarly, FMs required a lot of tuning by human artists, but provided a solid foundation for concept art and proof-of-concept user interfaces.
@@
Since production assistance is also close to traditional PCG pipelines, it is also understandably explored for games. Nasir and Togelius [115] used GPT-3 to generate levels for the Metavoidal (Yellow Lab Games, 2022) brawler game from a prompt that describes the level’s features (e.g. width and height) while a human curates and edits the results to ensure playability. This curated set of levels is then used for further fine-tuning, potentially automating the generative process. Kelly et al. [116] instead use GPT-4 to generate stories in natural language while abiding to logic constraints, assisting story writers. Instead of generating the final artifact directly, Kumaran et al. [117] use the LLM to parse game level parameters from the user request. The natural language input to the LLM contains information about both what the user is looking for and properties that the level of Future Worlds —a game-based learning exhibit about environmental sustainability [118] —should posses, such as difficulty, size, type of sustainability problem, and simulation goal metrics. A collection of candidate game levels are generated from a single user request. Each candidate is then evaluated by a RL agent, and the best game is presented to the user via the Unity game engine.
Figure 6: A screenshot of the LLMaker digital game content design assistant tool [119]. Image used with permission.
@@
Examples of procedural assistance are very few: LLMaker [119] (see Figure 6) is a chat-only level editor where the user can chat with the LLM and request changes to be made to the level and its content. Natural conversation with the LLM allows the designer to explore the current state of the level (for example, the user may ask “What are the enemies in the entrance room?”) or draw inspiration for the next changes to make (for example, by asking “What kind of enemies should I add to this corridor?”). When the user request a change to be made to the level, the LLM translates their request into a valid function call, ensuring content consistency and adherence to domain constraints [120]. Parameters for the function call that have not been specified by the user are generated by the LLM itself, biased by the existing level and overall user preferences. For example, if the user asks “Change the name of the banshee to Scary Ghost” the LLM will not only change the enemy’s name but also their description to reflect the change, even though this was not specified explicitly by the user (see Fig. 6) In addition the level and its content, LLMaker also employs Stable Diffusion models to generate the graphical assets for the artifacts, based on descriptions generated by the LLM itself. As noted above, so far research has mostly focused on either LLMs for conceptual assistance (putting significant onus on a human designer) or as production assistance (leveraging a human designer as curator). The conversational nature of LLMs, however, seems particularly well-suited for procedural assistance when designing content; we revisit this underexplored area of research in Section 4.

### v4

@@ 3.5 Analyst
So far, language models (both large and small) have been used for clustering player behaviors. Player2vec [69] trained the Longformer transformer architecture [70], with up to 121 million parameters, on a corpus of game events stored in JSON format which captured player interactions with a casual mobile game. Using dimensionality reduction methods on the latent vectors of the last layer of a pre-trained transformer, the authors discovered eight clusters with player traits that could be useful for market research, such as “lean-in casual economy aware” [69]. In such work, the strength of LLMs is their ability to process structured or unstructured data without the need for extensive data pre-processing. However, as with previous experiments in bottom-up clustering of players [71, 72], the resulting clusters need to be interpreted by expert game analysts (or the game’s designers) to derive some meaningful player types. In this form, current research points to LLMs acting as analyst assistants rather than independent analysts; we revisit future directions in this vein in Section 4.
Representing game logs as text interpreted by an LLM need not be the end-goal for analysis. The LLM representations of game logs can also be used to find common patterns that can be used for other purposes such as gameplay (or gameplay footage) similarity. Indicatively, Rasajski et al. [73] used LLMs on recorded gameplay action logs to establish action similarity. This similarity was in turn applied to train video encodings on gameplay footage of these recorded actions, in order to align the visual latent vectors to the action logs. While the work focused on the computer vision task of representation learning for gameplay footage, better (learned) representations of gameplay pixels would be more generalizable across games in downstream tasks. Such latent representations would be useful for general action recognition or general affect modeling [74]; we expand on these downstream tasks in Section 4.
@@ 3.6 Game Master
A Game Master (GM) in tabletop role-playing games (TTRPGs) is the person who creates the plot of a game, its characters, and narrative. GMs wear many hats during the course of the game session [75]; they prepare and adapt stories before sessions, guide gameplay during, and follow up with players afterward [76]. Digital games have mostly pre-scripted stories or level progressions and their players have a restricted range of affordances, compared to TTRPG players whose actions are only limited by their imagination. Similarly, the story told around the table can take any direction. Since human GMs mostly communicate about the gameworld, story, game state and action resolutions via natural language (although props such as maps, miniatures, hand-outs are also common), the potential of LLMs as a GM is often mentioned both in research circles and TTRPG discussion boards. LLMs as GMs also open the potential for solo play, while a TTRPG requires at least one player and a human GM.
@@
One of the first notable text adventures managed by a fine-tuned version of GPT-2 is AI Dungeon [77]. It is an online [^4] https://play.aidungeon.com/ interactive chat-based storytelling application where the player takes actions through language input alone. The LLM continues the story based on the player’s input, in the fashion of a human GM. The game has evolved since its creation to make use of more recent LLM models, which the player can choose from before starting a play session. Different gameworld settings are also offered, and players are also able to share the stories they create. Similar games have emerged online ever since [^5] https://koboldai.net, https://www.hiddendoor.co, and a freely available code repository, Kobold AI Client [^6] https://github.com/KoboldAI/KoboldAI-Client, allows a local or remote installation of a client for such LLM-run games. Some of these games also use Stable Diffusion text-to-image models [78] to generate visuals accompanying different parts of the narrative. More recent work [79] investigates how different characterizations of the GM and their way of presenting events to the players impacts the overall experience in TTRPGs. This study presented an online interface for a custom game based on a Dungeons & Dragons (TSR, 1974) fantasy setting. LLMs were ascribed different GM roles through a number of prompts, and human players participated in 45-minute games (split into 3 sub-sessions), with an overall positive response to the automated GM.
@@
In lieu of replacing a human GM, LLMs have also been employed as GM assistants. CALYPSO [80] is a set of tools running on a Discord server which the GM can query either to generate random encounters, brainstorm ideas, or alternatively chat with a fictional character in a Dungeons & Dragons setting. CALYPSO highlights that hallucinations of GPT-3 can have both positive effects when it generates plausible details not included in descriptions published in the original game manual (e.g. the shapes of creatures’ eyes) and also negative effects when the created details are outright incorrect (e.g. describing the wings of a canonically wingless creature). In addition, the model’s preconditioning to avoid racial bias was found to occasionally prevent it from generating racial details of fantasy creatures in the game. Other work used smaller GPT models to improvise in-game conversations [81] by monitoring and transcribing verbal exchanges between the GM and the players, and attempting to generate appropriate responses. This example was integrated into Shoelace [82], which is itself a GM assisting tool helping with content lookup by creating a node-based plan of the game narrative and encounters. The versatility of LLMs given their ability to rapidly process text input paves the way for their integration into the multitude of existing tools and aids for human GMs.
@@ 3.7 Game Mechanic
Another strength of LLMs is language synthesis. Huang and Sun [86] used GPT-3 to generate new words from the combination of two user-selected words. These are used to progress in a text-based game scenario, where the player wins by unlocking goal words. Similarly, language synthesis is leveraged by Infinite Craft [^8] https://neal.fun/infinite-craft/, an “alchemy” game, in which the player combines elements to produce new ones (see Fig. 3). In Infinite Craft, the player begins with a set of core elements (water, fire, wind and earth). But while the former have a set of interactions defined manually by the designer, Infinite Craft prompts Llama 2 [20] to imagine the product of the combination of these elements [87]. Judging from gameplay, it appears that for each distinct combination, Llama is prompted to produce the result only once, with the product stored in a database for future reference. Thus seemingly anything in the language model’s vocabulary might “emerge” from the combination of these elements, including all 50 states [^9] https://x.com/FeralFlex/status/1758332430136615298, “Dream” [^10] https://x.com/slutzsmp/status/1760394123243135169, and the fictional “Super Stonedosaurus Tacosaurus Rex” [^11] https://x.com/pvtspicy/status/1759316982984237139. On occasion, the model can choose to return one of the combined elements, or refuse to combine (e.g. very lengthy or complex) elements.
@@ 3.8 Automated Designer
Any PCG method that is trained on available content corpora fits under the Procedural Content Generation via Machine Learning (PCGML) paradigm [23]. Strictly speaking the original PCGML framework of 2018 did not consider LLMs; instead it relied on machine learning methods such as autoencoders and LSTMs. However, important challenges of PCGML remain when considering LLMs for PCG: notably, the reliance on high-quality, machine-readable datasets from human-authored levels. While some datasets exist for arcade game levels [88], for most games the content remains both unavailable and protected by intellectual property (IP) laws. We revisit this issue in Section 6. Prior work in PCG has demonstrated that tile-based game levels can be reliably generated with sequence-based prediction models (e.g. LSTMs) from a modest set of examples, by treating such levels as linear sequences of tile types in raster order [89, 90].
More recently, this approach has been extended to modern LLMs pre-trained on natural language (instead of sequence models trained from the ground up). Todd et al. [91] fine-tuned a GPT-2 model on a large dataset of Sokoban (Thinking Rabbit, 1982) levels and, at test time, sampled from the model to produce novel puzzles (see Figure 4). Interestingly, their results indicate that while the GPT-2 model struggles when the fine-tuning dataset is restricted in size, GPT-3 (and, presumably, larger models released since then) are better able to accommodate limited training sets.
@@
Using a similar approach, MarioGPT trains a GPT-2 model on a relatively small dataset of Super Mario Bros (Nintendo, 1985) levels [93]. MarioGPT overcomes the issue of data scarcity by using the initial dataset as the starting point for an evolutionary algorithm. Existing levels are selected and then sections of the level are mutated by sampling from the GPT model and then correcting the border between the re-generated section and the rest of the level with a similarly-trained BERT (i.e. bi-directional) model [2]. This approach produces a large and diverse set of playable levels, despite starting from less than 20 levels.
The above GPT-based level generation approaches also show the promise of incorporating natural language instructions to produce conditional level generators, either by prefixing game levels in the training dataset with desired level characteristics [91] or by embedding user instructions and allowing the model to attend to the embedding during generation [93]. A recent example of the latter is Cardistry [94], which used GPT-3.5 to transform a short personal narrative into a set of playing cards. Along with text information, the LLM creates prompts for DALL-E [95], completing the generation pipeline with playing card art.
User requests however are typically fuzzy and can be misinterpreted by the LLM. Hu et al. [56] first “refine” user requests for the generation of Minecraft structures using an LLM, adding knowledge specific to the domain (such as block palette and building dimensions). Then, a second LLM is employed to generate the description of the structure that can be interpreted and placed in the game.
While most applications of LLMs as automated designers involve some mapping from the LLM’s natural language output to other game-specific formats, LLMs can naturally be applied to games where the content is itself just natural language. In this vein, LLMs have been leveraged for the generation of New York Times Connections puzzles [96], which involve having a user group 16 words into 4 groups of 4 words each, with the challenge being for the user to infer what common semantic theme unites each group.
Figure 5: A screenshot of a CrawLLM game instance, generated for the theme of Ancient Egypt [97]. In CrawLLM, the LLM generates themes, stories, characters, and locations for a card-based dungeon crawler game while Stable Diffusion generates the visuals. Image used with permission.
@@
Finally, LLMs can generate new games by writing game code directly. For instance, the GAVEL system [99] combines evolution with LLMs to generate board games. In GAVEL, a language model fine-tuned on a dataset of board game programs written in the Ludii description language [100] acts as the mutation operator within the evolutionary loop.
@@ 3.9 Design Assistant
An AI for design assistance can provide several benefits to the creative process. Depending on the type of tool, type of AI, and type of creative process, the AI can minimize development time and cost, reduce human effort, support collaboration among members of a design team, or elicit a user’s creativity [101]. So far, in games, most of the AI-powered design assistant tools focus on autocompleting a human’s in-progress design [102] or providing many possible suggestions for the designer to consider [103, 104, 105, 106, 107, 108]. Ideally, we would want an LLM that can act as a human colleague that we can bounce off ideas to and collaborate with. Such LLMs are still beyond the current state of the art [109]. Existing tools implement co-creating [110] LLMs at different levels of control, which we can explore under the existing Co-Creative Framework for Interaction Design [111]. Focusing on the interaction over the artifact, the LLM can be of conceptual assistance, providing high-level guidance which is not game-ready. This would require that the designer adapts and curates the AI output in a way that fits their own vision and the constraints of the game. This implies that the LLM contributes very little in the actual artifact generation, merely providing new suggestions. By allowing the LLM to also refine or transform the artifact, we see a procedural assistance by the LLM. Interacting with the designer, the LLM can produce increasingly more final versions of the intended artifact. The LLM is expected to understand the context of the game for which the content is intended, in order to provide meaningful assistance. However, the LLM does not need to produce a final, playable artifact but could instead simply provide the next creative step for discussion with the designer [112]. Moreover, the designer is ultimately responsible for curating and adapting the generated content, as well as deciding when the co-creative process is completed [113, 114, 115]. Finally, if the LLM is allowed to directly create and alter the artifact based on user requirements, we say it provides production assistance. This is the closest level to PCG (see Section 3.8), but is different in that the designer remains in control and can refine their specification or reject a created artifact (versus an autonomous generator which directly sends content to the player). As expected, however, the AI operates in a much more constrained space in this scenario as it must account for all other game mechanics (the design of which are presumed finalized) and designer goals which are somehow encoded or presumed via learned designer models [116].
@@
One can argue that existing interfaces with LLMs and Large Multimodal Models (LMMs) act as design assistants. The designer provides their specifications and receives one (in LLMs) or multiple (in AI image generators) suggestions that they can further refine. Many creatives report using such interfaces for brainstorming and concept development [117], including game developers [118]. However, the applicability of LLMs as design assistants is somewhat limited, reverting only to conceptual assistance. Similarly, their potential for refining an existing idea (i.e. offering procedural assistance) is underexplored, as we discuss in Section 4.
@@
Conceptual assistance is thus the easiest for LLMs, and is the first case explored in games. Charity et al. [107] envision design assistance as a tool that combines the game description provided by the user with existing knowledge of similar games to suggest possible game features back to the user. The suggested features are fairly generic, few-word guidelines (e.g.: “learn new combat”) which would need extensive design effort and creativity to transform into an implementable and coherent game design. When asking an LLM for specific game features to implement in a digital game, players found them less compelling than human-designed ones [119]. The suggestions, however, were still useful for game designers, as they provided a different perspective that could kickstart their creation process. A thorough analysis of strengths and weaknesses of LLMs in this role can be drawn from Project AVA [120], a non-commercial digital game developed at Keywords Studios with the assistance of LLMs and LMMs for multiple aspects of the typical game development pipeline. LLMs are shown to help greatly in giving inspiration to the designers (albeit not being creative themselves), provide simple starting code for the game logic, and even assist developers by revealing errors in the code during development. LLMs however often fall short on anything more involved or requiring further domain knowledge, such as requesting code for specific game logic or understanding functionality in UI elements. Similarly, LMMs required a lot of tuning by human artists, but provided a solid foundation for concept art and proof-of-concept user interfaces.
@@
Since production assistance is also close to traditional PCG pipelines, it is also understandably explored for games. Nasir and Togelius [121] used GPT-3 to generate levels for the Metavoidal (Yellow Lab Games, 2022) brawler game from a prompt that describes the level’s features (e.g. width and height) while a human curates and edits the results to ensure playability. This curated set of levels is then used for further fine-tuning, potentially automating the generative process. Kelly et al. [122] instead use GPT-4 to generate stories in natural language while abiding to logic constraints, assisting story writers. Instead of generating the final artifact directly, Kumaran et al. [123] use the LLM to parse game level parameters from the user request. The natural language input to the LLM contains information about both what the user is looking for and properties that the level of Future Worlds—a game-based learning exhibit about environmental sustainability [124]—should posses, such as difficulty, size, type of sustainability problem, and simulation goal metrics. A collection of candidate game levels are generated from a single user request. Each candidate is then evaluated by a RL agent, and the best game is presented to the designer via the Unity game engine.
Figure 6: A screenshot of the LLMaker digital game content design assistant tool [125]. In LLMaker, a human designer makes requests in natural language to alter a dungeon level for a video game, editing both the level layout and the entities within. The LLM satisfies the designer requests by calling the appropriate function, ensuring that the content adheres to domain constraints. Image used with permission.
@@
Examples of procedural assistance are very few: LLMaker [125] (see Figure 6) is a chat-only level editor where the user can chat with the LLM and request changes to be made to the game level and its content. Natural conversation with the LLM allows the designer to explore the current state of the level (for example, the user may ask “What are the enemies in the entrance room?”) or draw inspiration for the next changes to make (for example, by asking “What kind of enemies should I add to this corridor?”). When the user requests a change to be made to the level, the LLM translates their request into a valid function call, ensuring content consistency and adherence to domain constraints [126]. Parameters for the function call that have not been specified by the user are generated by the LLM itself, biased by the existing level and overall user preferences. For example, if the user asks “Change the name of the banshee to Scary Ghost” the LLM will not only change the enemy’s name but also their description to reflect the change, even though this was not specified explicitly by the user (see Fig. 6) In addition to the level and its content, LLMaker employs Stable Diffusion models to generate the graphical assets for the artifacts, based on descriptions generated by the LLM itself. As noted above, so far research has mostly focused on either LLMs for conceptual assistance (putting significant onus on a human designer) or as production assistance (leveraging a human designer as curator). The conversational nature of LLMs, however, seems particularly well-suited for procedural assistance when designing content; we revisit this underexplored area of research in Section 4.